A 1.50kg object hangs motionless from a spring with a force constant of k=250N/m.how far is the spring stretched from it equilibrium length

1
Expert's answer
2021-11-29T11:46:53-0500

The weight of the object is:


"W = mg"

where "m = 1.50kg,g=9.81N\/kg". According to the Hook's law, the elongation of the spring is:


"x = \dfrac{W}{k}"

where "k=250N\/m". Thus, obtain:


"x = \dfrac{mg}{k} = \dfrac{1.50kg\cdot 9.81N/kg}{250N/m}\approx 0.0589m"

Answer. 0.0529m.